OpenStudy (anonymous):

A student is running a 3-kilometer race. He runs 1 kilometer every 2 minutes. Select the function that describes his distance from the finish line.

OpenStudy (anonymous):

f(x) = - 1 over 3x + 2 f(x) = - 1 over 2x + 3 f(x) = 1 over 2x + 3 f(x) = 1 over 3x + 2

OpenStudy (anonymous):

@phi?

OpenStudy (phi):

I guess they are saying let x be the time in minutes At the beginning of the race (at time x=0) he has to run 3 km if you replace x with 0 in the first choice \[ f(x) = -\frac{1}{3}x + 2\] what do you get ?

OpenStudy (anonymous):

You get 2. :)

OpenStudy (phi):

and that means the first equation means: at the start of the race he has to run 2 km But we know the race is 3 km. In other words, choice A is wrong.

OpenStudy (phi):

For the same reason, the choice D is also wrong.

OpenStudy (anonymous):

So is it B? :)

OpenStudy (phi):

It has to be B or C to figure out which, we use common sense. After running for 2 minutes, the distance from the finish line should be less (he is getting closer) in other words, if we replace x with 2, in equation B we get \[ f(x) = -\frac{1}{2}2 + 2 = -1+2 = 1\] in choice C we get \[ f(x) = \frac{1}{2}2 + 2= 1+2 =3 \]

OpenStudy (phi):

which equation gives a reasonable number for his distance from the finish line.

OpenStudy (anonymous):

C? Because he is running a 3 kilometer race? :)

OpenStudy (phi):

no. You want the function that describes his distance from the finish line in other words, an equation that tells you how far away he is from the finish line. after running (toward the line), he gets closer. the distance from the line gets less.

OpenStudy (anonymous):

Ohhh...I get it. So it is B :)

OpenStudy (phi):

yes

OpenStudy (anonymous):

Thank you thank you thank you! Now I get it. :)

OpenStudy (phi):

yw

OpenStudy (anonymous):

:D